If a and b are both positive and unequal, and logab + logba2 =3.

A. b= √a B. b= (√a2)2 C. b=a D. b=1/2 a


1
Expert's answer
2021-10-21T10:17:03-0400

If "a" and "b" are both positive and unequal, and "\log_ab + \log_ba^2 =3." It follows that "\log_ab + \frac{\log_aa^2}{\log_ab} =3," and hence "\frac{\log^2_ab +2\log_aa}{\log_ab} =3." Therefore, "\log^2_ab +2 =3\log_ab," which is equivalent to "\log^2_ab-3\log_ab +2 =0," and thus to "( \log_ab-1)(\log_ab -2) =0." We conclude that "\log_ab=1" or "\log_ab=2." It follows that "\log_ab=\log_aa" or "\log_ab=2\log_aa=\log_aa^2." Consequently, "b=a" or "b=a^2,\ a>0."

Taking into account that "a" and "b" must be both positive and unequal, we conclude that "b=a^2=((\sqrt{a})^2)^2" for "a>0."


Answer: B